# Do Language Models Use Their Depth Efficiently?

### v2

@@ Abstract
Modern LLMs are increasingly deep, and depth correlates with performance, albeit with diminishing returns. However, do these models use their depth efficiently? Do they compose more features to create higher-order computations that are impossible in shallow models, or do they merely spread the same kinds of computation out over more layers? To address these questions, we analyze the residual stream of the Llama 3.1 and Qwen 3 family of models. We find: First, comparing the output of the sublayers to the residual stream reveals that layers in the second half contribute much less than those in the first half, with a clear phase transition between the two halves. Second, skipping layers in the second half has a much smaller effect on future computations and output predictions. Third, for multihop tasks, we are unable to find evidence that models are using increased depth to compose subresults in examples involving many hops. Fourth, we seek to directly address whether deeper models are using their additional layers to perform new kinds of computation. To do this, we train linear maps from the residual stream of a shallow model to a deeper one. We find that layers with the same relative depth map best to each other, suggesting that the larger model simply spreads the same computations out over its many layers. All this evidence suggests that deeper models are not using their depth to learn new kinds of computation, but only using the greater depth to perform more fine-grained adjustments to the residual. This may help explain why increasing scale leads to diminishing returns for stacked Transformer architectures.
@@ 3 Experiments
Most of the experiments presented in the main paper are performed with Llama 3.1 70B [20], using NDIF and NNsight [21]. Unless noted otherwise, the results are computed on 10 random examples from GSM8K [22]. In bar plots, each bar starts from 0 (no stacking). The main results are also shown in the appendix on different models, including Llama and Qwen [23]. In Sec. 3.1, we measure how the layers and sublayers contribute to the residual stream. In Sec. 3.2, we use causal interventions to measure the effect of layers on downstream computations. In Sec. 3.3 we show that deeper or otherwise more complex computations do not influence the number of layers that have a causal effect on the prediction model. In Sec. 3.4 we train linear projections to find the correspondence between the layers of an independently trained shallow and deep Qwen model. Our exploration in Sec. 3.5 suggests that MoEUT [16] might use its layers more efficiently, especially when not modeling the question.111Our code is public: https://github.com/robertcsordas/llm_effective_depth
@@ 3.3 Do Deeper Problems Use Deeper Computation?
In order to more systematically verify if the models are using deeper computations of later operations, we measure the max influence of layers on later layers and the output on future positions, similarly to Fig. 3b and 13b. For each layer, we take the mean effect on all future layers in future tokens, thus reducing the maximum effect matrix to a vector with a dimension equal to the number of layers. For the separation point between the past and future tokens, we sample multiple positions from the answer. Additionally, we further reduce these “effect size” vectors by computing . Here, is the importance of layer computed by any of our previously defined metrics. We call the “depth score”. This score increases if the model uses the later layers more.
@@ 3.4 Do Deeper Models Do Novel Computation?
Are deeper models performing computation that is not present in the shallower ones because of a lack of layers, or is it “stretching out” the same kind of computation over more layers, performing smaller steps at a time? The first kind of effect would be preferable, because the deeper model is capable of combining more features and composing more subresults in theory. If such novel computation is present, it should be hard to predict from the activations of the shallow model, and the number of predictable layers should be close to the number of layers in the shallow model. In order to verify this, we take two pretrained models with different layer counts ( and , with ), and train different linear probes to map every point in the residual of the shallower model to each representation in the deeper model . This requires that the two models use the same tokenizer, and, for reliable results, they should be trained independently, instead of being distilled from each other. Because of GPU memory limitations, we decided to use the Qwen 2.5 series of models because of their more modest size compared to the Llama models. Concretely, we train a linear map for each pair of layers of Qwen 2.5 1.5B and 14B. We measure the relative prediction error , where is the linear map from layer of the small model to of the big model. We do this for each layer pair and plot it in Fig. 9. Although some ranges of layers seem to be easier to predict than others, a clear diagonal pattern is visible. This shows that the big model is more likely a “stretched out” version of the shallow model, rather than one that does entirely new computations.
@@ 3.5 Is the Pretraining Objective or the Model Responsible for using Fixed Depth?
We use the 244M parameter baseline and MoEUT models from the paper [16], without modifications. We perform the residual erasure intervention on four examples (see Appendix Sec. C.5), and display the most important results in Fig. 9 (with details in Figs. 31 and 32 in the Appendix). We can clearly see that the models that were trained to not model the uncertain question use more of their layers in processing the answer, probably because they do not have to spend their capacity on modeling the high-entropy probability distribution of the unpredictable questions. Surprisingly, MoEUT successfully achieves this even when learning to model the question, although the effect is more pronounced without it. This confirms the advantage of the shared-layer models. Interestingly, while all models have good interpolation performance, their extrapolation capability differs drastically. For example, MoEUT on the “Mul/Div Multiple Longer” split has an accuracy of 36% if modeling the question is enabled, and 63% if it is not. The difference for standard Transformers is less dramatic (41 vs. 48%).
Interestingly, all models trained from scratch show increased depth with deeper computation steps. The effect is more pronounced with the MoEUT models, especially if the question is not learned. This is in contrast to what we found for most examples when fine-tuning a Llama model (App. C.5), confirming that fine-tuning might not be enough to change the pretrained model’s behavior fundamentally.
@@ 4 Related Work
For additional related work, please refer to Appendix E.
@@ Appendix B Model Depth as a Factor Shaping Performance
Appendix C Results on Other Models
@@
C.1 How do the Layers Interact With the Residual Stream?
We show the absolute and relative contributions of the sublayers to the residual stream for Llama 3.1 8b in Fig. 14, and for the Qwen 3 series of models in Fig 15. We show cosine similarities of the sublayer’s contributions and the residual stream for all Llama and Qwen models that we tested in Fig. 16. The results are similar to our findings in Sec. 3.1.
@@ C.1 How do the Layers Interact With the Residual Stream?
C.2 How do the Layers Influence Downstream Computations?
Here, we show the effect of individual layers on later layers in future timesteps and on future token outputs for multiple models. Fig 17 shows the effect on Llama 3.1 8B and 70B, while Fig. 18 shows the Qwen 3 series of models, and Fig. 19 shows the instruction tuned Llama 3.1 70B. It can be seen that instruction tuning has no influence on the model’s behavior. In the Qwen 3 series of models, the effect seems to be less pronounced, but still present. The findings agree with our discussion in Sec. 3.2.
We show the local layer interactions for other Llama models in Fig. 20 and for the Qwen 3 models in Fig. 21.
Additional Logitlens results are shown for the other models in Fig. 22.
@@ C.2 How do the Layers Influence Downstream Computations?
C.3 Do Deeper Problems Use Deeper Computation?
Additional residual erasure experiments are shown for Llama 3.1 70B in Fig. 23 and for the Qwen 3 series of models in Fig. 24. Findings for the 70B Llama models are identical to those discussed in Sec. 3.3. Qwen models use more layers, but they also seem to use a fixed number of layers independently of the computation depth, indicating that they are not building on subresults from previous computation steps. This is also consistent with the findings of Sec 3.3.
C.4 Do Deeper Models Do Novel Computation?
@@
C.5 Does Finetuning Cause the Model to Use Deeper Computations?
@@
Appendix D Details on the DeepMind Math Training
@@
Appendix E Extended Related Work
@@
Appendix F Hardware Resources
@@
Full-finetuning Llama 3.1 3B on the DeepMind Math Dataset (Sec. C.5) was done on 4 Nvidia H200 GPUs for 10 hours.

### v3

@@ Abstract
Modern LLMs are increasingly deep, and depth correlates with performance, albeit with diminishing returns. However, do these models use their depth efficiently? Do they compose more features to create higher-order computations that are impossible in shallow models, or do they merely spread the same kinds of computation out over more layers? To address these questions, we analyze the residual stream of the Llama 3.1, Qwen 3, and OLMo 2 family of models. We find: First, comparing the output of the sublayers to the residual stream reveals that layers in the second half contribute much less than those in the first half, with a clear phase transition between the two halves. Second, skipping layers in the second half has a much smaller effect on future computations and output predictions. Third, for multihop tasks, we are unable to find evidence that models are using increased depth to compose subresults in examples involving many hops. Fourth, we seek to directly address whether deeper models are using their additional layers to perform new kinds of computation. To do this, we train linear maps from the residual stream of a shallow model to a deeper one. We find that layers with the same relative depth map best to each other, suggesting that the larger model simply spreads the same computations out over its many layers. All this evidence suggests that deeper models are not using their depth to learn new kinds of computation, but only using the greater depth to perform more fine-grained adjustments to the residual. This may help explain why increasing scale leads to diminishing returns for stacked Transformer architectures.
@@ 3 Experiments
Most of the experiments presented in the main paper are performed with Llama 3.1 70B [20], using NDIF and NNsight [21]. Unless noted otherwise, the results are computed on 10 random examples from GSM8K [22]. In bar plots, each bar starts from 0 (no stacking). The main results are also shown in the appendix on different models, including Llama, Qwen [23], and OLMo 2 [24]. In Sec. 3.1, we measure how the layers and sublayers contribute to the residual stream. In Sec. 3.2, we use causal interventions to measure the effect of layers on downstream computations. In Sec. 3.3 we show that deeper or otherwise more complex computations do not influence the number of layers that have a causal effect on the prediction model. In Sec. 3.4 we train linear projections to find the correspondence between the layers of an independently trained shallow and deep Qwen model. Our exploration in Sec. 3.5 suggests that MoEUT [16] might use its layers more efficiently, especially when not modeling the question.111Our code is public: https://github.com/robertcsordas/llm_effective_depth
@@ 3.3 Do Deeper Problems Use Deeper Computation?
In order to more systematically verify if the models are using deeper computations of later operations, we measure the max influence of layers on later layers and the output on future positions, similarly to Fig. 3b and 17b. For each layer, we take the mean effect on all future layers in future tokens, thus reducing the maximum effect matrix to a vector with a dimension equal to the number of layers. For the separation point between the past and future tokens, we sample multiple positions from the answer. Additionally, we further reduce these “effect size” vectors by computing . Here, is the importance of layer computed by any of our previously defined metrics. We call the “depth score”. This score increases if the model uses the later layers more.
@@ 3.4 Do Deeper Models Do Novel Computation?
Are deeper models performing computation that is not present in the shallower ones because of a lack of layers, or is it “stretching out” the same kind of computation over more layers, performing smaller steps at a time? The first kind of effect would be preferable, because the deeper model is capable of combining more features and composing more subresults in theory. If such novel computation is present, it should be hard to predict from the activations of the shallow model, and the number of predictable layers should be close to the number of layers in the shallow model. In order to verify this, we take two pretrained models with different layer counts ( and , with ), and train different linear probes to map every point in the residual of the shallower model to each representation in the deeper model . This requires that the two models use the same tokenizer, and, for reliable results, they should be trained independently, instead of being distilled from each other. Because of GPU memory limitations, we decided to use the Qwen 2.5 [33] series of models because of their more modest size compared to the Llama models. Concretely, we train a linear map for each pair of layers of Qwen 2.5 1.5B and 14B. We measure the relative prediction error , where is the linear map from layer of the small model to of the big model. We do this for each layer pair and plot it in Fig. 9. Although some ranges of layers seem to be easier to predict than others, a clear diagonal pattern is visible. This shows that the big model is more likely a “stretched out” version of the shallow model, rather than one that does entirely new computations.
@@ 3.5 Is the Pretraining Objective or the Model Responsible for using Fixed Depth?
We use the 244M parameter baseline and MoEUT models from the paper [16], without modifications. We perform the residual erasure intervention on four examples (see Appendix Sec. D.5), and display the most important results in Fig. 9 (with details in Figs. 42 and 43 in the Appendix). We can clearly see that the models that were trained to not model the uncertain question use more of their layers in processing the answer, probably because they do not have to spend their capacity on modeling the high-entropy probability distribution of the unpredictable questions. Surprisingly, MoEUT successfully achieves this even when learning to model the question, although the effect is more pronounced without it. This confirms the advantage of the shared-layer models. Interestingly, while all models have good interpolation performance, their extrapolation capability differs drastically. For example, MoEUT on the “Mul/Div Multiple Longer” split has an accuracy of 36% if modeling the question is enabled, and 63% if it is not. The difference for standard Transformers is less dramatic (41 vs. 48%).
Interestingly, all models trained from scratch show increased depth with deeper computation steps. The effect is more pronounced with the MoEUT models, especially if the question is not learned. This is in contrast to what we found for most examples when fine-tuning a Llama model (App. D.5), confirming that fine-tuning might not be enough to change the pretrained model’s behavior fundamentally.
@@ 4 Related Work
For additional related work, please refer to Appendix F.
@@ Appendix B Model Depth as a Factor Shaping Performance
Appendix C Robustness Analysis
To justify our choice of a relatively low number of samples (10-20, depending on the experiment) and a single dataset (GSM8k), we present results for more samples and for a different dataset (Math). We show the influence of layers and sublayers on the residual stream, equivalent to Fig. 2, in Fig. 10. Furthermore, we show the layer skipping effects, equivalent to Fig. 3, in Fig. 11, and the Logitlens-based output similarity, corresponding to Fig. 4, in Fig. 12. These analyses show that our conclusions are robust both with respect to the dataset choice and the number of examples.
To show the variability between individual examples, we show 4 examples contributing to Fig. 3b in Fig. 13. Their variability justifies our choice to take the maximum over them to quantify the overall importance of the layers.
Appendix D Results on Other Models
@@
D.1 How do the Layers Interact With the Residual Stream?
We show the absolute and relative contributions of the sublayers to the residual stream for Llama 3.1 8b in Fig. 18, for the Qwen 3 series of models in Fig 19 and for OLMo 2 models in Fig 20. We show cosine similarities of the sublayer’s contributions and the residual stream for all Llama and Qwen models that we tested in Fig. 21 and for the OLMo models in Fig. 22. The results are similar to our findings in Sec. 3.1.
@@ D.1 How do the Layers Interact With the Residual Stream?
D.2 How do the Layers Influence Downstream Computations?
Here, we show the effect of individual layers on later layers in future timesteps and on future token outputs for multiple models. Fig 23 shows the effect on Llama 3.1 8B and 70B, while Fig. 24 shows the Qwen 3 series of models, Fig. 25 shows the instruction-tuned Llama 3.1 70B, and Fig. 26 shows the OLMo 2 models. It can be seen that instruction tuning has no influence on the model’s behavior. In the Qwen 3 series of models, the effect seems to be less pronounced, but still present. OLMo 2 behaves very similarly to Qwen 3. The findings agree with our discussion in Sec. 3.2.
We show the local layer interactions for other Llama models in Fig. 27, for the Qwen 3 models in Fig. 28 and for OLMo 2 in Fig. 29.
@@ D.2 How do the Layers Influence Downstream Computations?
D.3 Do Deeper Problems Use Deeper Computation?
Additional residual erasure experiments are shown for Llama 3.1 70B in Fig. 32, for the Qwen 3 series of models in Fig. 33 and for OLMo 2 in Fig. 34. Findings for the 70B Llama models are identical to those discussed in Sec. 3.3. Qwen models use more layers, but they also seem to use a fixed number of layers independently of the computation depth, indicating that they are not building on subresults from previous computation steps. OLMo seems to have weaker depth dependence than the other models. All findings are consistent with Sec 3.3.
D.4 Do Deeper Models Do Novel Computation?
@@
D.5 Does Finetuning Cause the Model to Use Deeper Computations?
@@
Appendix E Details on the DeepMind Math Training
@@
Appendix F Extended Related Work
@@
Previously, some methods were proposed that might help increase the effectiveness of the feature building in deeper layers. For example, Kim et al. [47] show that their method increases the angular distance between the representations of different layers, indicating that it might use deeper layers more efficiently. Li et al. [48] mix post-layernorms in early layers and pre-layernorms in later layers to improve gradient propagation. However, we are not aware of freely available popular large language models using these configurations in order to verify their effectiveness. However, we tested the OLMo 2 series of models with a different layernorm structure, resulting in similar conclusions as the standard pre-layernorm models.
The collapse of importance of attention in the later layers could be related to massive activations [49] that have been shown to collapse the attention to a few discrete tokens, potentially reducing their effectiveness. It might also partially explain the success of selective attenetion pruning [50].
@@
Appendix G Hardware Resources
@@
Full-finetuning Llama 3.1 3B on the DeepMind Math Dataset (Sec. D.5) was done on 4 Nvidia H200 GPUs for 10 hours.